# Circle bundles with PSC over some four manifolds

### v2

@@ Abstract.
We construct infinitely many examples of four manifolds with macroscopic dimension $4$ equipped with circle bundles whose total spaces admit metrics of positive scalar curvature. Further, we verify that these bundles have macroscopic dimension at most $3$. In particular, we answer a question of Gromov on the existence of circle bundles over enlargeable manifolds whose total spaces admit metrics of positive scalar curvature. Our constructions are based on techniques from symplectic geometry.
The second author is supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500.
@@ 1.1. Background and motivation
Remark.
Note that for a compact metric space $X$, $\dim_{mc}(X)=0$. Therefore, when one refers to the macroscopic dimension of a closed manifold $M$, one typically means the macroscopic dimension of the universal cover $\widetilde{M}$.
@@
While Conjecture 1 arises from the simple observation that existence of a PSC metric is guaranteed given the presence of an $S^{2}$ factor in the manifold, it is also natural to ask about the possible interaction between existence of PSC metrics and having an $S^{1}$ factor. This has been widely investigated. For instance, PSC manifolds with a trivial $S^{1}$ factor (i.e., $M\times S^{1}$) are amenable to the Schoen-Yau conformal descent argument [SY79], as they have a non-trivial codimension-$1$ homology class given by the Poincaré dual of the $S^{1}$ factor. Such $S^{1}$ factors are also useful in quantitative geometric problems related to scalar curvature. For example, building on an observation of Fisher-Colbrie and Schoen [FCS80], Gromov and Lawson introduced the toric symmetrization procedure [GL83] which was used to estimate the Lipschitz constant of a non-zero degree map from a PSC manifold. More recently, Gromov [Gro18] observed that toric symmetrization can also be used to prove width estimates for Riemannian bands. On the same theme, Rosenberg [Ros07, Conjecture 1.24] formulated the following conjecture regarding the relationship between PSC metrics and trivial $S^{1}$ factors.
Conjecture 2 (Rosenberg).
Let $M$ be a closed connected $n$-dimensional manifold. Then $M\times S^{1}$ admits a PSC metric iff $M$ admits a PSC metric.
However, the situation is less clear when the $S^{1}$ factor arises as a fiber of a non-trivial circle bundle. The key difficulty in studying existence of PSC metrics on non-trivial circle bundles is that, in this case, the Schoen-Yau conformal descent argument does not apply in a straightforward way. [^1] Indeed, the integral homology class of the $S^{1}$ fiber in nontrivial circle bundles is always torsion, with order given by the divisibility of the Euler class of the bundle. Therefore, its Poincaré dual is zero. In this direction, Gromov posed the following question [Gro18, pg. 661]:
@@
In the negative direction we point out the recent comprehensive work [He23] on PSC metrics on non-trivial circle bundles due to He. In particular, he showed that circle bundles with homologically non-trivial fibers do not admit PSC metrics if the base space is a $1$-enlargeable2221111-enlargeable means that the maps to $S^{n}$ in the definition of enlargeability are of degree $1$. manifold of dimension $n\leq 6$ [He23, Proposition 4.12].
The main goal of this article is to study and exhibit examples of interactions between the existence of a positive scalar curvature metric and macroscopic dimension on non-trivial circle bundles over some $4$-manifolds. In particular, we give a positive answer to Question 1. We construct these examples using techniques from symplectic geometry. The inspiration for our construction came from the following remark by Gromov [Gro23, pg. 8]:
The emergent picture of spaces with $Sc.curv\geq 0$, where topology and geometry are intimately intertwined, is reminiscent of the symplectic geometry, but the former has not reached yet the maturity of the latter.
Granting this dictum of similarity between scalar curvature geometry and symplectic geometry, one is led to believe that the analogue of the Schoen-Yau descent for positive scalar curvature manifolds, in the symplectic case must be Donaldson’s symplectic divisor theorem [Don96], as it can be viewed as a codimension-$2$ descent-type result for symplectic manifolds. This analogy is strengthened when seen in conjunction with a recent result of Cecchini, Räde and Zeidler [CRZ23] which establishes a certain codimension-$2$ descent for positive scalar curvature manifolds of dimension $7$. In fact, we show that circle bundles over Donaldson divisors in symplecic $6$-manifolds fit neatly into the framework of applicability of [CRZ23].
@@ 1.2. Main results
In this article, we construct infinitely many examples of $4$-manifolds $Z$ with arbitrarily large Euler number $|\chi(Z)|$, arising as Donaldson divisors (see Theorem 1) in symplectic $6$-manifolds, such that $Z$ does not admit a PSC metric and macroscopic dimension of $\widetilde{Z}$ is exactly $4$. Nevertheless, we construct circle bundles over these $Z$,
@@
[…]
@@
To the best of our knowledge examples exhibiting a drop in macroscopic dimension on twisting with $S^{1}$ have not appeared before in literature. On a similar theme, we were informed by Guth [Gut24] that a drop in the Urysohn width upon taking a double cover, i.e. an $S^{0}$ bundle, was observed by Balitskiy [Bal21, Theorem 2.4.4, pg. 36].
@@
    The total space $E_{Z}$ admits a metric of positive scalar curvature.
@@
    The macroscopic dimension of the total space $E_{Z}$ is at most 3.
@@
    If $b_{2}^{+}(M)>1$, then $Z$ does not admit a positive scalar curvature metric.
@@
    If $M$ admits a positive degree map to an aspherical manifold, then $Z$ does not admit a positive scalar curvature metric.
  5. (5)
    If $M$ is a symplectically aspherical [^4] See Definition 3. In particular, a symplectic manifold with $\pi_{2}=0$ (e.g. $T^{4}$) is symplectically aspherical. manifold and $\pi_{1}(M)$ is amenable, then the macroscopic dimension of $Z$ is exactly $4$.
We emphasize that Conclusions $(1)$ and $(2)$ make no assumptions either on the fundamental groups of $M$ and $E_{Z}$, or on the (non-)existence of spin or almost spin structures on the universal cover of $E_{Z}$. In particular, Conclusion $(2)$ does not follow from any existing results on Gromov’s conjecture (Conjecture 1). On the other hand, Conclusions $(3)$, $(4)$ and $(5)$ rest on Lemma 5, where we will show that there is a positive degree map from $Z$ to $M$. Then Conclusion $(3)$ follows from a theorem of Taubes [Tau94], Conclusion $(4)$ follows from the work of Chodosh, Li and Liokumovich [CLL23] and Conclusion $(5)$ follows from combining Proposition 6 with a result of Dranishnikov [Dra11].
By combining Theorem A with a Euler characteristic computation for Donaldson divisors carried out in Proposition 4, we obtain the following uniqueness corollary:
Corollary B.
There exists infinitely many $4$-manifolds $Z_{k}$ and circle bundles $E_{k}$ over $Z_{k}$ such that $|\chi(Z_{k})|\sim O(k^{3})$, and
  1. (1)
    $\dim_{mc}Z_{k}=4$ but $\dim_{mc}E_{k}=3$,
  2. (2)
    $Z_{k}$ is not PSC and $E_{k}$ is PSC.
As a further corollary of Theorem A, we will construct examples sought after in Gromov’s Question 1.
Corollary C.
There exists enlargeable closed $4$-manifolds $Z$ and nontrivial circle bundles $E\to Z$ such that $E$ admits a positive scalar curvature metric.
@@ 1.3. Organisation
In Section 2, we define the necessary symplectic preliminaries and make some observations regarding Donaldson divisors. In Section 3, we assemble other lemmas and propositions that we need to prove our main results. In Section 4, we prove Theorem A and Corollary C as well as provide explicit examples that are used towards proofs of Corollary B. The reader may first read Section 4, and refer to Sections 2 and 3, for the proofs of the results used in Section 4 as they arise.
@@ Acknowledgements
The authors would like to thank Shihang He for several helpful comments on the first version of this paper. They would also like to thank Alexey Balitskiy, Alexander Dranishnikov, and Larry Guth for answering their questions and providing helpful remarks. The first author wishes to thank his advisor Yannick Sire for support and encouragement. The second author wishes to thank his advisor Mahan Mj, as well as his friends and colleagues Ritwik Chakraborty, and Sekh Kiran Ajij for several helpful conversations.
@@ 2. Donaldson divisors: definition and properties
Remark.
Any symplectic manifold $(X,\eta)$ admits a symplectic form $\omega$ which is integral. To see this, choose a collection of closed $2$-forms on $X$ representing a $\mathbf{Q}$-basis of the vector space $H^{2}(X;\mathbf{Q})\subset H^{2}(X;\mathbf{R})$, and expresses the class $[\eta]\in H^{2}(X;\mathbf{R})$ as a $\mathbf{R}$-linear combination of the chosen basis elements. Approximate the real coefficients by rational numbers and then multiply by an integer to clear out the denominators. Let $\omega$ be the resulting closed $2$-form. Then $\omega$ is an integral symplectic form on $X$. Indeed, one only needs to verify that $\omega$ is non-degenerate. But it is so, since non-degeneracy is both an open condition and is scale-invariant.
@@ 2.1. Complement of Donaldson divisors
By the construction of $Z\subset X$ summarized above, there is an asymptotically holomorphic section $s:X\to L^{\otimes k}$ such that $Z=s^{-1}(\mathbf{0})$. Let $\psi:X\setminus Z\to\mathbf{R}$ be defined by $\psi:=-\log|s|^{2}$. During the course of the proof of [Don96, Proposition 39], the author shows that $\psi$ is a Morse function on $X\setminus Z$ with critical points having index at most $n$. Therefore, by Morse theory, $X\setminus Z$ deformation retracts to an $n$-dimensional simplicial complex. ∎
@@
    If $n\geq 3$, the inclusion $\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$.
@@
For Part $(2)$, observe that the Morse function $\psi$ constructed in Lemma 2.2 gives a handlebody decomposition of $W$ by handles of index at most $n$. By reversing the handlebody decomposition, we get that $W$ can be constructed from $\partial W$ by attaching handles of index at least $n$. If $n\geq 3$, then we conclude that all these handles are of index at least $3$. Since $\pi_{1}$ is unchanged under attachements of handles of index bigger than $2$, we conclude $\partial W\hookrightarrow W$ induces an isomorphism in $\pi_{1}$. ∎
2.2. Topology of Donaldson divisors in dimension $6$
In the next proposition, we show that Donaldson divisors of a $6$-dimensional symplectic manifold $(X^{6},\omega)$ with integral symplectic form $\omega$, realizing the homology classes Poincaré dual to $k[\omega]$, are topologically distinct for distinct values of $k\gg 1$. We do this by calculating their Euler characteristic.
@@ 2.2. Topology of Donaldson divisors in dimension $6$
Let $Z_{k}\subset(X^{6},\omega)$ denote a Donaldson divisor with $[Z_{k}]=\mathrm{PD}(k[\omega])$. Then, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|\sim O(k^{3})$ is monotonically increasing in $k$. In particular, for $k,l\gg 1$ sufficiently large, $Z_{k}\cong Z_{l}$ if and only if $k=l$.
@@
Substituting (2.3) in (2.2), we obtain
$c_{2}(TZ_{k})$ $=j^{*}c_{2}(TX)-j^{*}c_{1}(TX)\smile c_{1}(L^{\otimes k}|_{Z_{k}})+k\cdot j^{*}[\omega]\smile c_{1}(L^{\otimes k}|_{Z_{k}})$
$=j^{*}c_{2}(TX)-k\cdot j^{*}(c_{1}(TX)\smile[\omega])+k^{2}\cdot j^{*}[\omega]^{\smile 2}$ (2.4)
Since the top Chern class is the Euler class, evaluating the cohomology classes appearing in either sides of (2.4) on $[Z_{k}]=k\cdot\mathrm{PD}[\omega]$ we get:
$\chi(Z_{k})$ $=\int_{Z_{k}}c_{2}(TZ_{k})$
$=k^{3}\int_{M}[\omega]^{\smile 3}-k^{2}\int_{M}c_{1}(TX)\smile[\omega]^{\smile 2}+k\int_{M}c_{2}(TX)\smile[\omega]$
Note that $c_{1}(TX),c_{2}(TX),[\omega]$ are all independent of $k$. Hence, $\chi(Z_{k})$ is a cubic polynomial in $k$ with coefficients independent of $k$, and a nonzero leading coefficient. Therefore, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|$ is monotonically increasing in $k$, as required. ∎
Remark.
The fact that $\chi(Z_{k})\sim O(k^{3})$ in $X^{6}$ should be put in context of the following standard observation: Let $X=\mathbf{CP}^{3}$ and $V_{k}\subset\mathbf{CP}^{3}$ be a degree $k$ complex hypersurface, then $\chi(V_{k})=k^{3}-4k^{2}+6k$.
2.3. Donaldson divisors of $M\times S^{2}$
Let $(M^{4},\eta_{0})$ be an integral symplectic $4$-manifold. We equip the manifold $M\times S^{2}$ with the integral symplectic form $\omega=\eta_{0}\oplus\omega_{0}$, where $\omega_{0}$ denotes the standard area form on $S^{2}$. We now make the following observation, which will be crucial to the proof of Parts (4) and (5) of Theorem A.

### v3

# Circle bundles with PSC over large manifolds

@@ Abstract.
We construct infinitely many examples of macroscopically large manifolds of dimension $m\geq 4$ equipped with circle bundles whose total spaces admit metrics of positive scalar curvature and have macroscopic dimension at most $\lceil m/2\rceil+1$. In particular, we answer a question of Gromov on the existence of circle bundles over enlargeable manifolds whose total spaces admit metrics of positive scalar curvature, in all dimensions. Our constructions are based on techniques from symplectic geometry.
@@ 1.1. Background and motivation
Remark 1.
Note that for a compact metric space $X$, $\dim_{mc}(X)=0$. Therefore, when one refers to the macroscopic dimension of a closed manifold $M$, one typically means the macroscopic dimension of the universal cover $\widetilde{M}$, equipped with a metric obtained from lifting some metric on $M$. As any two metrics on a closed manifold are Lipschitz equivalent, this quantity is independent of the chosen metric on $M$.
@@
While Conjecture 1 arises from the simple observation that existence of a PSC metric is guaranteed given the presence of an $S^{2}$ factor in the manifold, it is also natural to ask about the possible interaction between existence of PSC metrics and having an $S^{1}$ factor. This has been widely investigated. For instance, PSC manifolds with a trivial $S^{1}$ factor (i.e., $M\times S^{1}$) are amenable to the Schoen-Yau conformal descent argument [SY79], as they have a non-trivial codimension-$1$ homology class given by the Poincaré dual of the $S^{1}$ factor. In fact, this was one of the key ideas in Schoen and Yau’s proof that for $n\leq 7$, the torus $T^{n}=S^{1}\times\cdots\times S^{1}$ does not admit a PSC metric. Thereafter, Gromov and Lawson [GL83] proved this for all $n$. An important concept in their proof was the notion of enlargeability. We now recall the definition of enlargeable manifolds from [GL83].
Definition 3 (Enlargeability).
The basic example of an enlargeable manifold is the unit circle $S^{1}\cong\mathbf{R}/2\pi\mathbf{Z}$. Indeed, the $n$-sheeted cover of $\mathbf{R}/2\pi\mathbf{Z}$ is $\mathbf{R}/2\pi n\mathbf{Z}$. The map $\mathbf{R}/2\pi n\mathbf{Z}\to\mathbf{R}/2\pi\mathbf{Z}$ given as scaling by $1/n$ is a $1/n$-contracting map, for any $n\geq 1$. Also, $S^{1}$ is certainly a spin manifold. It is straightforward to see that
  1. (1)
    Product of enlargeable manifolds is enlargeable,
  2. (2)
Note that if $M$ is enlargeable, then $M\times S^{1}$ is also enlargeable as the product of two enlargeable manifolds is enlargeable. However, the situation is less clear when the $S^{1}$ factor arises as a fiber of a non-trivial circle bundle. In this direction, Gromov posed the following question [Gro18, pg. 661]:
@@
In the negative direction we point out the recent comprehensive work due to He [He23] on PSC metrics on non-trivial circle bundles. In particular, He shows that circle bundles with homologically non-trivial fibers do not admit PSC metrics if the base space is a $1$-enlargeable1111111-enlargeable means that the maps to $S^{n}$ in the definition of enlargeability are of degree $1$. manifold of dimension $n\leq 6$ and the circle fibers are homologically non-trivial in the total space [He23, Proposition 4.12].
The main goal of this article is to study and exhibit examples of interactions between the existence of a positive scalar curvature metric and macroscopic dimension on non-trivial circle bundles over manifolds of dimension $n\geq 4$. In particular, we give a positive answer to Question 1. We construct these examples using techniques from symplectic geometry. The inspiration for our construction came from a remark due to Gromov [Gro23, pg. 8]:
“The emergent picture of spaces with $Sc.curv\geq 0$, where topology and geometry are intimately intertwined, is reminiscent of the symplectic geometry, but the former has not reached yet the maturity of the latter.”
Granting this dictum of similarity between scalar curvature geometry and symplectic geometry, one is led to believe that the analogue of the Schoen-Yau descent for positive scalar curvature manifolds, in the symplectic case must be Donaldson’s symplectic divisor theorem [Don96], as it can be viewed as a codimension-$2$ descent-type result for symplectic manifolds. This analogy serves as the starting point of our constructions.
@@ 1.2. Main results
For $n\geq 2$, in this article, we construct infinitely many examples of $2n$-manifolds $Z$, arising as Donaldson divisors (see, Theorem 1) in symplectic $(2n+2)$-manifolds, such that $Z$ does not admit a PSC metric and macroscopic dimension of $\widetilde{Z}$ is exactly $2n$. Nevertheless, we construct circle bundles over these $Z$,
@@
such that the $(2n+1)$-dimensional total space $E_{Z}$ admits a PSC metric. An intriguing feature of these examples is that $\dim_{mc}(\widetilde{E_{Z}})\leq n+1$. Thus, even though $E_{Z}$ is obtained from $Z$ by taking a twisted product with $S^{1}$, strangely the macroscopic dimension drops by half for $E_{Z}$. For $n=2$, note that this is sharply in consonance with Gromov’s macroscopic dimension conjecture (Conjecture 1), i.e., $\dim_{mc}(\widetilde{E_{Z}})\leq\dim E_{Z}-2$. We show this using the natural deformation retraction from $\widetilde{E_{Z}}$ to an $(n+1)$-dimensional simplicial complex arising as the spine of the complement of the Donaldson divisor $Z$ in the ambient symplectic $(2n+2)$-manifold.
@@
To the best of our knowledge, examples exhibiting a drop in macroscopic dimension upon twisting with $S^{1}$ have not appeared before in literature. On a similar theme, we were informed by Guth [Gut24] that a decrease in Urysohn width upon taking a double cover, i.e. an $S^{0}$ bundle, was observed by Balitskiy [Bal21, Theorem 2.4.4, pg. 36].
@@
    The total space $E_{Z}^{2n+1}$ admits a metric of positive scalar curvature.
@@
    The macroscopic dimension of the total space $E_{Z}$ is at most $n+1$.
@@
    Let $(M^{2n+2},\omega)$ be a symplectically aspherical [^2] See Definition 4. In particular, any symplectic manifold with $\pi_{2}=0$ (e.g. $T^{2n}$) is symplectically aspherical. manifold with an amenable fundamental group. Then the macroscopic dimension of $Z^{2n}$ is exactly $2n$.
@@
    Let $M^{2n+2}=N^{2n}\times S_{g}$ and $\omega=\omega_{N}\oplus\omega_{0}$, where $\omega_{0}$ is an integral area form on $S_{g}$. If $N$ is enlargeable and $k$ is even, then $Z$ is also enlargeable. In particular, $Z$ does not admit a metric of positive scalar curvature.
Additionally, varying the integer $k\gg 1$ produces an infinite collection of non-homeomorphic manifolds $Z^{2n}$ satisfying the above.
We emphasize that the Conclusions $(1)$ and $(2)$ does not rely on any assumptions either on the fundamental groups of $M$ and $E_{Z}$, or on the (non-)existence of spin or almost spin structures on the universal cover of $E_{Z}$. In particular, Conclusion $(2)$ does not follow from any existing results on Gromov’s conjecture (Conjecture 1).
On the other hand, Conclusions $(3)$ and $(4)$ rest on Proposition 5, where we will show that there is a positive degree map from $Z$ to $M$. Then, Conclusion $(3)$ follows from combining Proposition 7 with a result of Dranishnikov [Dra11], and Conclusion $(4)$ follows from the fact that a manifold admitting a positive degree map to an enlargeable manifold is also enlargeable, provided it admits a finite cover that is spin. In this case, it follows from the classical result of Gromov and Lawson [GL80] that $Z$ does not admit a PSC metric. Finally, it follows from Proposition 4 that varying the parameter $k$ produces an infinite collection of such manifolds $Z$ satisfying the conclusions of Theorem A.
Conclusion $(4)$ of Theorem A gives the examples sought after in Gromov’s question (Question 1) in all even dimension $\geq 4$. The case of odd dimensions $>4$ follows upon considering the bundle $p\times\mathrm{id}:E_{Z}\times S^{1}\to Z\times S^{1}$ obtained by taking product of the bundle $p:E_{Z}\to Z$ with $S^{1}$, and observing that $Z\times S^{1}$ is enlargeable if $Z$ is enlargeable.
Hanke [Han25] communicated to us a proof showing that total spaces of circle bundles over $3$-manifolds which are either enlargeable or macroscopic dimension $3$, do not admit positive scalar curvature metrics. We record the proof in Section 3.2.
Theorem B.
Let $M$ be a closed connected $3$-manifold that is either enlargeable or has macroscopic dimension $3$. Let $p:E\to Z$ be a circle bundle. Then the total space $E$ does not admit a metric of positive scalar curvature.
Remark 2.
In dimension $2$, the only closed connected manifolds which are enlargeable or have macroscopic dimension $2$ are surfaces of genus $g\geq 1$, or their non-orientable quotients. Since these are aspherical, total spaces of circle bundles over such surfaces are also aspherical. By the classification of PSC $3$-manifolds, these can not admit positive scalar curvature metrics.
On the other hand, in dimension $1$, the total spaces of such circle bundles are either diffeomorphic to a torus or a Klein bottle. These can not admit a metric of positive scalar curvature either, by the Gauss-Bonnet theorem.
Theorem A together with Theorem B provides a complete answer to Gromov’s question (Question 1) for closed manifolds in all dimensions: if $\dim Y\geq 4$, there are infinitely many examples of circle bundles $\underline{X}\to Y$ where $Y$ is enlargeable and $\underline{X}$ admits a metric with $Sc>0$. On the other hand, if $\dim Y\leq 3$, then there are no such examples.
@@ 1.3. Organisation
In Section 2, we define the necessary symplectic preliminaries and make some observations regarding Donaldson divisors. In Section 3, we prove Theorem A and Theorem B. The reader may first read Section 3, and refer to Section 2 for the proofs of the results used in Section 3 as they arise. In Section 4, we discuss several examples of manifolds satisfying the conclusions of Theorem A, each illustrating different possible properties of the manifolds produces by our constructions.
@@ Acknowledgements
The authors would like to thank Shihang He for several helpful comments on the first version of this paper. They would also like to thank Alexey Balitskiy, Alexander Dranishnikov, and Larry Guth for answering their questions and providing several helpful remarks. The authors are also indebted to Bernhard Hanke for his invaluable input on an earlier version of the paper which helped greatly simplify the proofs, as well as for communicating a proof of Theorem B. The second author wishes to thank his advisor Mahan Mj for his interest, as well as his friends and colleagues Ritwik Chakraborty and Sekh Kiran Ajij for several helpful conversations. The second author is supported by the Department of Atomic Energy, Government of India, under project no.12-R&D-TFR-5.01-0500.
@@ 2. Donaldson divisors: definition and properties
Remark 3.
Any symplectic manifold $(X,\eta)$ admits a symplectic form $\omega$ which is integral. To see this, choose a collection of closed $2$-forms on $X$ representing a basis of the vector space $H^{2}(X;\mathbf{R})$, and express the class $\eta$ as an $\mathbf{R}$-linear combination of the chosen closed $2$-forms. Since $H^{2}(X;\mathbf{Q})\subset H^{2}(X;\mathbf{R})$ is dense, we may approximate the real coefficients by rational numbers and then multiply by an integer to clear out the denominators. Let $\omega$ be the resulting closed $2$-form. Then $[\omega]\in H^{2}(X;\mathbf{Z})$. Since non-degeneracy is both an open condition and is scale-invariant, $\omega$ is a non-degenerate $2$-form. Therefore, $\omega$ is an integral symplectic form, as desired.
@@ 2.1. Complement of Donaldson divisors
By the construction of $Z\subset X$ summarized above, there is an asymptotically holomorphic section $s:X\to L^{\otimes k}$ such that $Z=s^{-1}(\mathbf{0})$. Let $\psi:X\setminus Z\to\mathbf{R}$ be defined by $\psi:=-\log|s|^{2}$. During the course of the proof of [Don96, Proposition 39], it is shown that $\psi$ is a Morse function on $X\setminus Z$ with critical points having index at most $n$. Therefore, by Morse theory, $X\setminus Z$ deformation retracts to an $n$-dimensional simplicial complex. ∎
@@
    If $n\geq 3$, the inclusion $\partial W\hookrightarrow W$ induces an isomorphism on $\pi_{1}$.
@@
For Part $(2)$, observe that the Morse function $\psi$ constructed in Lemma 2.2 gives a handlebody decomposition of $W$ by handles of index at most $n$. By reversing the handlebody decomposition, we get that $W$ can be constructed from $\partial W$ by attaching handles of index at least $n$. If $n\geq 3$, then we conclude that all these handles are of index at least $3$. Since $\pi_{1}$ is unchanged under attachements of handles of index bigger than $2$, we conclude $\partial W\hookrightarrow W$ induces an isomorphism on $\pi_{1}$. ∎
2.2. Topology of Donaldson divisors
In the next proposition, we show that Donaldson divisors of a symplectic manifold $(X^{2n},\omega)$ with integral symplectic form $\omega$, realizing the homology classes Poincaré dual to $k[\omega]$, are topologically distinct for distinct values of $k\gg 1$. We do this by calculating their Euler characteristic.
@@ 2.2. Topology of Donaldson divisors
Let $Z_{k}\subset(X^{2n},\omega)$ denote a Donaldson divisor with $[Z_{k}]=\mathrm{PD}(k[\omega])$. Then, for sufficiently large values of $k\gg 1$, $|\chi(Z_{k})|\sim O(k^{n})$ is monotonically increasing in $k$. In particular, for $k,l\gg 1$ sufficiently large, $Z_{k}\cong Z_{l}$ if and only if $k=l$.
@@
Solving for the system of equations given by (2.3) and (2.2), we obtain
$c_{n-1}(TZ_{k})=\sum_{i=0}^{n-1}(-)^{i}k^{i}\cdot j^{*}c_{n-1-i}(TX)\smile[\omega]^{\smile i},$ (2.4)
where $c_{0}(TX)=1$ by definition. Since the top Chern class is the Euler class, evaluating the cohomology classes appearing in either sides of (2.4) on $[Z_{k}]=k\cdot\mathrm{PD}[\omega]$ we get:
$\chi(Z_{k})=\int_{Z_{k}}c_{n-1}(TZ_{k})$ $=\int_{X}c_{n-1}(TZ_{k})\smile k[\omega]$
$=\sum_{i=1}^{n}(-1)^{i-1}k^{i}\int_{X}c_{n-1}(TX)\smile[\omega]^{i}$
Note that $[\omega]$ and $c_{i}(TX)$, $0\leq i\leq n$ are all independent of $k$. Hence, $\chi(Z_{k})$ is a polynomial in $k$ with degree $n+1$, with coefficients independent of $k$, and the leading coefficient being nonzero. Therefore, for $k\gg 1$, $|\chi(Z_{k})|$ is monotonically increasing in $k$, as required. ∎
Remark 4.
The fact $|\chi(Z_{k})|\sim O(k^{n})$ should be put in context of the following standard observation: let $X=\mathbf{CP}^{n}$ and $V_{k}\subset\mathbf{CP}^{n}$ be a degree $k$ complex hypersurface. Then,
$\chi(V_{k})=((1-k)^{n+1}-1)k^{-1}+n+1.$
2.3. Donaldson divisors of $M\times S_{g}$
Let $S_{g}$ be a closed surface with genus $g\geq 0$. Let $(M^{2n},\eta_{0})$ be an integral symplectic manifold. We equip the manifold $M\times S_{g}$ with the integral symplectic form $\omega=\eta_{0}\oplus\omega_{0}$, where $\omega_{0}$ denotes a positive, integral area form on $S_{g}$. We now make the following observation, which will be crucial to the proof of Part (4) of Theorem A.